OpenStudy (anonymous):

Lt (1/n)^(1/ln n)=? n->infty

OpenStudy (anonymous):

\[Y=\lim_{n \rightarrow \infty} (1/n)^{1/ \ln n}\] Taking logarithm \[\ln Y=\lim_{n \rightarrow \infty} (1/\ln n)\times(\ln (1/n))\] \[\ln Y=\lim_{n \rightarrow \infty} (1/\ln n)\times[-\ln (n)]\] \[\ln Y=\lim_{n \rightarrow \infty} (-1)\] \[\ln Y=(-1)\] \[Y=e^{-1}\]